OpenStudy (anonymous):

increasing a number by 5 and multiplying the result by 7 is the same as multiplying the number by -5 and subtracting 1 from the result. what is the number?

OpenStudy (anonymous):

7(x+5)= -5x-1 7x+35= -5x-1 7x+5x=-1-35 12x=-36 x=-3

OpenStudy (anonymous):

oh my gosh, thankyou!